OpenStudy (anonymous):

Find a cubic function with the given zeros. sqrt7, -sqrt7, -4

OpenStudy (anonymous):

this is easy. you want an equation which spits out 0 when x= sqrt7, -sqrt7 and -4. i.e the roots of the equation are x=sqrt7 x=-sqrt7 x=-4 this implies that x-sqrt7=0 ..........1 x+sqrt7=0 .........2 x+4=0 ..........3 Now you get the equation by just multiplying 1,2 and 3. i.e. \((x-\sqrt{7})(x+\sqrt{7})(x+4)=0.\)

OpenStudy (anonymous):

you can simplify that if you want.

OpenStudy (anonymous):

would it be f(x) = x^3 + 4x^2 - 7x - 28?

OpenStudy (anonymous):

idk you tell me...all you have to do is multiply those terms. I haven't simplified it.